Homework Problem Set #2 A W10x22 used as a column under axial compression. Assume that it is unbraced unless specified otherwise. Use A36 steel. For each column length and end conditions given, calculate the design capacity, pePa (25 points) (a) L = 8 ft, pinned ends (b) L-8 ft, fixed ends (c) L = 16 ft, pinned ends (d) L = 16 ft, fixed ends (e) L- 16 ft, pinned ends, braced at the midspan in the weak axis direction.
Answer #1

solatsonGwhwI0x 22 10(8x12:12.18く200 (ok) KL - 91 4:귀 |트 133.68 36 E3-2 ㅈ2 (29000) = 54.94 ks. (36/S444)36 = 27-365 ksi Te mơ

EA,-,6.48(649)冫106-56kips ANS De»ign th 106 S6kibs (b〉 L-8ft , find ends ·65 (8x12. 1.33 6.92)

(36/30-oi F32-08(4-49) - 208 012 kips 九 em - #xs 14-286 (6:49), 124.84 kips

Maxnmum 144.364200(ok) = 91 1323 -r2 (29ooo 13-74 KL ,2.045 (6.49〉 78.172 kíps ェ LRED. design stengte AS : kad enls Since 93.

43-84)L 36 22,645 ks R, -FO A 22 . ơ 4 5 6.49 14 6.966 kips tl is = o.FF0.-ος (22.64s)ニ13. 587 ks. 13.5 8 4 6.49S ェ 88.18 ki